OpenStudy (anonymous):

The probability that a student takes a history class and a sociology class is 0.051. The probability that a student takes a history class is 0.32. What is the probability that a student takes a sociology class given that the student is taking a history class? 0.051 0.159 0.269 0.32

OpenStudy (anonymous):

Use Bayes' rule: \[P(A|B) = \frac{P(A \textrm{ and } B)}{P(B)}\]

OpenStudy (anonymous):

So, it's \[\frac{0.051}{0.32} = 0.159\]

OpenStudy (anonymous):

thank you! do you mind helping me with another question on probability?

OpenStudy (anonymous):

Yeah, sure

OpenStudy (anonymous):

A political candidate wanted to understand what issues were important to their voters. They asked 1000 voters, "Which issue is the most important: crime, city planning, or health care?" The results of the survey are shown in the two-way table below: Crime City Planning Health Care Male 118 239 458 Female 418 317 449 What is the probability that a person chosen at random from this survey is concerned about crime given that they are male? Round your answer to the nearest tenth. 5.9% 14.5% 22.0% 26.8%

OpenStudy (anonymous):

So, you could use Bayes' rule, but it's easier just to disregard the female statistics and only look at the male line. The probability is 118/(118+239+458) = 0.145 = 14.5%.